An institute reported that 70?% of its members indicate that lack of ethical culture within financial firms has contributed most to the lack of trust in the financial industry. Suppose that you select a sample of 100 institute members.


The probability is 60?% that the sample percentage will be contained above __?% and below __?%.
Answer #1

Normal approximation to the binomial calculation:

Standard error of the mean = SEM = ?x(N-x)/N3 = 0.046

? = (1-CL)/2 = 0.200

Standard normal deviate for ? = Z? = 0.842

Proportion of positive results = P = x/N = 0.700

Lower bound = P - (Z?*SEM) = 0.661

Upper bound = P + (Z?*SEM) = 0.739

The probability is 60% that the sample percentage will be contained above _66.1% and below 73.9%.?